OpenStudy (anonymous):

Can you tell me if this is right. (y^2 -6y+8)(y+2) (y^2-6y+8)y +(y^2 -6y +8)2 y(y^2-6y+8) +2(y^2-6y+8) y*y^2 - y*-6y +y*8 + 2*y^2 - 2*-6y +2*8 y^3-6y^2+8y +2y^2+12y+16 y^3-4y^2+20y+16 I am just not sure if I put my minuses and pluses in the right spot. Please explain if it is wrong.

OpenStudy (anonymous):

y^3-4y^2-4y+16 it is -12y + 8y

myininaya (myininaya):

oh wait you wrote extra negative signs in the step before

OpenStudy (anonymous):

Can you walk me through it from the beginning, or just we can work through it together.

myininaya (myininaya):

y^3+y*(-6y)+8y+2y^2-6y*2+16 y^3-6y^2+2y^2-12y+8y+16 y^3-4y^2-4y+16

OpenStudy (anonymous):

Then lets start at: (y^2-6y+8)(y+2)

OpenStudy (anonymous):

so it is not y*y^2?

myininaya (myininaya):

the first term is y*y^2=y^3

myininaya (myininaya):

y*(y^2-6y+8)=y^3-6y^2+8y 2*(y^2-6y+8)= 2y^2-12y+16 ------------------- y^3-4y^2-4y+16

OpenStudy (anonymous):

Oh ok, so my first sentence is right. y*y^2-y*6y+y*8+2*y^2 -2*-6y+2*8 So if there is anything wrong with this line let me know, is there?

myininaya (myininaya):

that is right

OpenStudy (anonymous):

So the answer is y^3-4y^2-4y+16